OpenStudy (vanz):

divide using synthetic division: (4x^3+7x^2-62x+15) / (4x-1)

OpenStudy (amistre64):

4x-1 = 0 when x = ?

OpenStudy (vanz):

??? :O

OpenStudy (amistre64):

3 2 1 0 <-- make sure all the parts are there; they are 4x^3 +7x^2 -62x +15 0 <-- add this row ---------------------- r ) <-- multiply this row ^ use the root

OpenStudy (amistre64):

the root is when 4x-1 = 0 ...

OpenStudy (amistre64):

im gonna say when x = 1/4 we get a zero ... 4x^3 +7x^2 -62x +15 0 1 2 -15 ---------------------- 1/4 ) 4 8 -60 [0], remainder

OpenStudy (mertsj):